OpenStudy (anonymous):

Solve: 2(x + 1) = −4

OpenStudy (anonymous):

2x + 2 = -4 2x = -4-2= -6 x = -3

OpenStudy (anonymous):

1) Factor 2(x+1) into 2x+2 (That is ,carry and multiply the 2 to both terms). 2x+2=-4 Now, substract the 2 from the left side and you get: 2x=-6 Divide the -6 by 2 to solve for x. x=-3!!